I know the answer to part b is supposed to be 3.77 x 10^-4 but I don't understand how.

1. A loop of wire (R- 10cm) with a counterclockwise current (I 30mA) is in a uniform magnetic field (B 0.4) directed to the right. The plane defined by the loop is parallel to the field. a (1p) If the loop is free to rotate, the loop will rotate such that the dot moves out of the page. ii. rotate such that the dot moves into the page. iii. not rotate at all b. (%pt) Calculate the magnitude of the initial torque on the loop (include units).
Answer #1

Solution:

b ) Torque on the loop = I A B

where I = 30 mA = 0.030 A

Magnetic field = B = 0.4 T

Area = A = pi R^2 = pi (0.10)^2 = 0.0314 m^2

Torque = (0.030)(0.0314)(0.4) = 3.77 x10^-4 N m